PRODUCT COST FLOWS. Use the following information to answer all parts of the question.

Materials Inventory
15*
?   25
10**
Work-in Process Inventory
10*
?   35
28
14
42**
Finished Goods Inventory
30* 55
?  
10**
Wages Payable
50
28
Manufacturing Overhead
12 14
2**

* Beginning Balance

** Ending Balance

Required

a. Cost of goods completed and transferred

?

b. Total manufacturing costs

?

c. Cost of goods manufactured

?

d. Cost of goods sold

?

e. Direct labor used

?
f. Overhead applied

?

g. Overhead incurred

?

h. Raw materials purchased

?

i. Raw materials used

?

j. Over-applied overhead



k. Under-applied overhead
Answer #1
Direct Materials
Beginning Raw Materials Inventory a 15
Add: Purchases of raw materials b 10
Deduct: Ending Raw Materials Inventory c 10
Direct Materials used in production d = a + b – c 15
Direct Labor e 42
Manufacturing Overhead f 1212
Total Manufacturing costs g = d + e + f 1269
Add: Beginning WIP Inventory h 30
Deduct: Ending WIP Inventory i 42
Cost of Goods Manufactured for the Year j = g + h – i 1257
Add: Beginning Finished Goods Inventory k 30
Deduct: Ending Finished Goods Inventory l 10
Cost of Goods Sold m = j + k – l

1277

answers

a) 1257

b) 1269

c) 1257

d) 1277

e) 28

f) 1212

g) 1214

h) =25-10 = available - initial inventory= 10

I) 15+10-10 =15 initial inventory + purchases - final inventory

j) 0

k) 2 it left to aply the diffence between 1214 incurred -1212 applied